Let the two vectors x & y and the matrix z be defined as follows 1.2 2.2 4.1 x-| 2.21, y-| 1.51,2-12.1-3.2 1.9 3.1 1.2 3.2 0.35 Write a script in Matlab and save it as .m file with name HW19_2. The script will execute the following tasks 1 Enter the vectors x &y and the Matrix z into the script. 2- Evaluate L2 lx2 3- Evaluate L1xl1 4- Evaluate Linf- l 5- Evaluate the dot product N-(x,y) 6- evaluated the angle in degrees between the vectors x & y 7- Evaluate M2 lzll2 8- Evaluate M1=111 9 Evaluate Minf zl
Answer #1

the matlab code for the problem is as follows:

clc; close all; clear all; x - [1;2.2;-1.35] y-[-2.4;1.5;0.35] z [-1.2 2.2 4.1; 2.1-3.2 1.9,3.1 1.2 3.2] L2 norm(x,2) L1 = norm(x,1) Linf = norm(x, inf) CosTheta -dot(x,y)/(norm(x)*norm(y)); ThetaInDegreesacosd (CosTheta); M2norm(z,2) M1norm(z,1) Minf = norm(z, inf)

the output from the command window after running the above code is